140624 LML Observations

Kona and Peter met at the LML site at 10:30 am.  No fog was collected, although the doors had opened.  We removed the strand frames and washed with DI at the sink.  We rinsed the collector body and drip tray/tube assembly in place with copious amounts of DI water.  Reassembled and reset for sampling tonight.

Leave a Reply

Your email address will not be published. Required fields are marked *